OpenStudy (anonymous):

A body has a KE E . If the mass of the body increases 25 times , Calculate its velocity , The kinetic energy remains the same .

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

Try...It By ur Own..)

OpenStudy (anonymous):

i dont know , give a hint !

OpenStudy (anonymous):

It is Almost Same as....The earlier one which i did.... Go through that Prob...U will get this

OpenStudy (anonymous):

I didn't get the question , what 25 times ?

OpenStudy (anonymous):

The mass Get multiplied...by 25

OpenStudy (anonymous):

yeah okay , ty !

OpenStudy (anonymous):

E = 1/2 * m * v1^2 E = 1/2 * 25 m * v2^2

OpenStudy (anonymous):

Can....u Do it Nw......u have to Find V2

OpenStudy (anonymous):

Go By the same way i did in previous Question..if u have Doubts ask

OpenStudy (anonymous):

okay , i have another question that says 4 is doubled , does that mean 4*2 or 4*4 ?

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

it means 4 is multiplied by 2

OpenStudy (anonymous):

okay ty yahoo :)

OpenStudy (anonymous):

welcome..)